OpenStudy (anonymous):

3/2x-15=-12 3/2(x+8)=6x

OpenStudy (anonymous):

so u want x price ???

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

\[\frac{3}{2x} -15 = -12\] \[\frac{3}{2(x+8)} = 6x\]

OpenStudy (anonymous):

it's (2) ok. or u want proof

OpenStudy (anonymous):

i really just need help with the second one, i keep trying and i do it wrong

OpenStudy (anonymous):

Is the second equation looks as I typed..?

OpenStudy (anonymous):

maybe it's mis written !

OpenStudy (anonymous):

no thats how it is

OpenStudy (anonymous):

Then distribute x and multiply to remove the denominator so you would get a quadratic equation.. :)